Derive a formula for Summation n i=1 i3 by using a telescoping sum with terms f(i)=i2

1
Expert's answer
2021-12-14T11:37:14-0500

i⁴ - (i-1)⁴ = i⁴ - (i⁴ - 4i³ + 6i² - 4i + 1)

=> i⁴ - (i-1)⁴ = i⁴ - i⁴ + 4i³ - 6i² + 4i - 1

=> i⁴ - (i-1)⁴ = 4i³ - 6i² + 4i - 1

Taking sum over i = 1 to n

"\sum_{i=1}^{n}" {i⁴ - (i-1)⁴ }= "\sum_{i=1}^{n}"{4i³- 6i²+4i - 1}

=> n⁴ - 0⁴ = "\sum_{i=1}^{n}4i³-\sum_{i=1}^{n} 6i² +\sum_{i=1}^{n} 4i - \sum_{i=1}^{n} 1"

=> n⁴ = "4\sum_{i=1}^{n}i³-6\sum_{i=1}^{n} i² +4\sum_{i=1}^{n} i - \sum_{i=1}^{n} 1"

=> n⁴ = "4\sum_{i=1}^{n}i³-" 6."\frac{n(n+1)(2n+1)}{6}" + "4.\frac{n(n+1)}{2}-n"

=> n⁴ = "4\sum_{i=1}^{n}i³" - n(n+1)(2n+1) + 2n(n+1) - n

=> "4\sum_{i=1}^{n}i³" = n⁴ + n(n+1)(2n+1) - 2n(n+1) + n

=> "4\sum_{i=1}^{n}i³" = n⁴ + 2n³ + 3n² + n - 2n² - 2n + n

=> "4\sum_{i=1}^{n}i³" = n⁴ + 2n³ + n²

=> "4\sum_{i=1}^{n}i³" = (n² + n)²

=> "4\sum_{i=1}^{n}i³" = {n(n+1)}²

=> "\sum_{i=1}^{n}i³" = "\frac{ {n²(n+1)}²}{4}"

=> "\sum_{i=1}^{n}i³ =" "[\frac{n(n+1)}{2}]²"